the area of a rectangle is x^2+6x+8 square units. if the length is x+4 units, find an algebraic expression for the perimeter

Answer by jim_thompson5910(35256)

x^2+6x+8 factors to (x+4)(x+2)

The length is x+4
The width is x+2

Perimeter:

P = 2*(L+W)
P = 2*(x+4+x+2)
P = 2*(2x+6)
P = 4x+12

The expression for the perimeter is 4x+12

------------------------------------------------------------------------------------------------------------------------

If you need more one-on-one help, email me at jim_thompson5910@hotmail.com. You can ask me a few more questions for free, but afterwards, I would charge you ($2 a problem to have steps shown or $1 a problem for answer only).

Alternatively, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Any amount is greatly appreciated as it helps me a lot. This donation is to support free tutoring. Thank you.

Jim
------------------------------------------------------------------------------------------------------------------------